Solve the linear system by graphing.
2x+y= -8
-x + y=4
Please help. I don't understand could you explain it in a very easy way please. Thanks

Answer by only4christ2(12)

Put each of the equations in y = mx + b form
2x + y = -8 becomes y = -2x - 8
-x + y = 4 becomes y = x + 4
Graph each on the same coordinate plane, using the slope and the y-intercept. Where the two lines intersect, that is where the solution is.
If you have a graphing calculator, you can graph them there for the solution as well.